OpenStudy (anonymous):

help with math review list all possible real roots, then solve each of the polynomial equations over the complex numbers. 8. x^3-7x^2-x+7=0 9. x^4+x^3-6x^2-14x-12=0

OpenStudy (anonymous):

8. x= 1, -1, 7 x^3 -7x^2-x=7 9. x= -2, 3, -1+ i, -1-i x^4 + x^3 -6x^2 -14x=12

OpenStudy (anonymous):

thanks